OpenStudy (anonymous):

Rearrange the equation 3y-6x+9=0 into the form y=____

OpenStudy (luigi0210):

Start by getting y by itself.

OpenStudy (anonymous):

y=2x-3

OpenStudy (anonymous):

3y-6x=-9 so then 3y=-9+6x ibut you want to single out the y, so it will be y=-9+6x over 3

OpenStudy (anonymous):

y=2x-3 is the simpliest form that you can reserve... if you do 3y= -9+6x or 3y=(6x-9) or y= (6x-9)/3---------------------(A) or y= 6x/3 - 9/3 or y= 2x-3------------------------(B) you stopped at (A) but i resolved and made it even simplier to (B)..

OpenStudy (anonymous):

both answer would be correct..

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

Put x= something. This time as an example, lets take x=0. When x=0, in (A) y=-3 When x=0, in(B) y=-3 so both the equations derieved are actually true and same the difference is that one is more simplified than the other.